OpenStudy (anonymous):

find the variation constant and an equation of variation when y varies inversely as x and 7=7 when x =11

OpenStudy (anonymous):

y=k/x k=yx

OpenStudy (anonymous):

y = k /x when y =7 (I thing that's what you meant) , 7 = k / 11 k = 77 y =k /x y = 77/x

OpenStudy (anonymous):

should be when y =7 when x=11

OpenStudy (anonymous):

You should better start doing it on your own. :]